OpenStudy (anonymous):

a plumber finished three jobs on tuesday the first two only cost the owner the 45$ trip fee because they took very little time to complete. For the third job, the plumber charged the trip fee plus 6 times his hourly rate.If the plumber received a total of 303$ for the day , what is the hourly rate ?

OpenStudy (anonymous):

who would even understand that??? and how much was the 2nd job? are u suposed to subtract 45 from 303 then divide by 6?

OpenStudy (anonymous):

thats the question im confused too !

OpenStudy (anonymous):

what grade r u in, @nicole06 i would just take 303-45= __ <=lets call this n than take n and multiply it by 6 and see wat u get. k?

OpenStudy (anonymous):

Total = job1 + job2 + job3 $303 = 45 + 45 + 45 + 6* r @nicole06 Can you finish it up?

OpenStudy (anonymous):

" he first two only cost the owner the 45$ trip fee"

OpenStudy (anonymous):

yeah, sorry. i missed that.

OpenStudy (anonymous):

@nicole06 .....???